the sarple mean to be uithia 0.25 of the poptioa nces with probabi lite - 935500?
Answer #1

z value at 93.85% = 1.8699

ME = 0. 25

ME = z *(s/sqrt(n))
0.25 = 1.8699 *(2.02/sqrt(n))
n = (1.8699/2.02/0.25)^2
n = 228